Hi, In class I was taught how to change an equation to function form here is an example y-x=1 or 8x+2y=-4, could you teach me how to change an equation to a function form?
Thanks

Answer by stanbon(75887)

y-x=1 or 8x+2y=-4, could you teach me how to change
an equation to a function form?
-----
Linear Form looks like y = mx + b
-----
Your 1st equation is y-x = 1
---
Solve it for "y"::
Add x to both sides to get:
y = x + 1
It is now in linear form.
------------------------
Your 2nd equation is 8x + 2y = -4
Solve for "y"::
Divide both sides of the equation by 2 to get:
4x + y = -2
---
Subtract 4x from both sides to get:
y = -4x -2
It is now in linear form.
====================
Cheers,
Stan H.
------------------